OpenStudy (anonymous):

Use the discriminate to determine the number of and type of solutions. -3m2 - 7m = 6

OpenStudy (anonymous):

probably easiest to start with \[3m^2+7m+6=0\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then compute \[b^2-4ac\] with \[a=,b=7,c=6\] and see what you get

OpenStudy (anonymous):

oops i meant \[b^2-4ac\] with \[a=3,b=7,c=6\]

OpenStudy (anonymous):

3m2+7m+6=0 compare with ax2 +bx+c a=3,b=7 and c=6 d=b^2-4ac d=-23 so it has complex roots

OpenStudy (anonymous):

two different complex solution as d<0

OpenStudy (anonymous):

okay, Thank you everbody. :)

OpenStudy (anonymous):

ur welcome